OpenStudy (anonymous):

Someone help! Convert standard form: y=x^2-2x-3 to vertex and intercept form. Convert vertex form: y-5=-2(x+1)^2 to standard and intercept form. Convert y=(x+2)(x-3) to standard and vertex form.

OpenStudy (anonymous):

Hmm give me a second to remember! what math are you in?

OpenStudy (anonymous):

Pre-calc/trig

OpenStudy (anonymous):

haha helping someone who is more advanced in math than me!! I'm only in Geometry, but i will do my best to help you out!!! i'll do some research and help ya!! do you think you can help me out?

OpenStudy (anonymous):

Ok thanks! I can try. What is your question?

OpenStudy (anonymous):

I'll post it right now!

OpenStudy (anonymous):

Okay i think i know how to do it

OpenStudy (anonymous):

Tell me if this makes any sense

OpenStudy (anonymous):

standard form to vertex form. standard Y=ax^2+bx+c Vertex y=a(x-h)^2+k vertex: (h,K) y=x^2-2x-3 Going to add 3 to both sides. y+3=x^2-2x ( )( )x^2-2x+ find C (x )(x )x^2-2x+

OpenStudy (anonymous):

Okay so i kinda go stuck on that part

OpenStudy (anonymous):

When I did it I got y=1(x-2)^2 -4, but I'm not sure if that's right

OpenStudy (anonymous):

well what did you get for your two binomials?

OpenStudy (anonymous):

The (x-1)^2

OpenStudy (anonymous):

Well to me that looks right. I mean, I'm not in this math class yet haha but to me when i tried to work the rest out it seemed right

OpenStudy (anonymous):

Ok thanks!

OpenStudy (anonymous):

no problem!